### v1

# Fast Scrambling in Hyperbolic Ising Model

Abstract
We investigate many-body chaos and scrambling in the Hyperbolic Ising model, a mixed-field Ising model living in the background of $AdS_{2}$. The effect of the curvature is captured by site-dependent couplings obtained from the $AdS_{2}$ metric applied to a flat nearest neighbor spin chain. We show that this model with only local site-dependent nearest neighbor interactions is maximally chaotic, can be classified as a fast scrambler, and saturates the Maldacena-Shenker-Stanford (MSS) bound on chaos for a certain set of parameters, thus making this model one of the few if not the only example where such fast scrambling behavior has been seen without all-to-all or long-range interactions. Moreover, the modest resources needed to simulate this model make it an ideal test-bed for studying scrambling and chaos on quantum computers.
I Introduction
Quantum many-body chaos and scrambling have attracted a great deal of attention in recent years, from more theoretical endeavors like studying the scrambling dynamics of black holes using their dual descriptions Kobrin et al. (2021); Fu and Sachdev (2016); Asaduzzaman et al. (2024a); Brown et al. (2023); Schuster et al. (2022); Cotler et al. (2017); Joshi et al. (2022) to more practical applications like investigating scrambling in chemical reactions to quantum cryptography Chamon et al. (2024, 2022) and many more.
In almost all of these studies, the Sachdev-Ye-Kitaev (SYK) model has been the center of attention as the quintessential example of a model that exhibits fast scrambling and quantum chaos Sachdev and Ye (1993); noa (a, b); Maldacena and Stanford (2016). However, due to the all-to-all and random interactions in this model, any kind of classical and quantum simulation can be very challenging and often requires various truncations and approximations Xu et al. (2020); García-García et al. (2021); Cao et al. (2020); Sünderhauf et al. (2019). In this letter, we advocate for a new model that is also maximally chaotic and exhibits fast scrambling but with only local site-dependent interactions.  [^1] There are other examples of fast scrambling models besides the SYK model but to the best of our knowledge they all require longer-range interactions than nearest neighbor interactions to be present in the model Hashizume et al. (2021); Li et al. (2020); Lashkari et al. (2013); Belyansky et al. (2020); Yin and Lucas (2020); Bentsen et al. (2019).
A wide variety of tests have been developed to diagnose and classify quantum chaos, including spectral form factors Kos et al. (2018); Liu (2018); Bertini et al. (2018), level statistics Berry and Tabor (1977); Bohigas et al. (1984), different definitions of complexityRoberts and Yoshida (2017); Balasubramanian et al. (2022), Loschmidt echos Peres (1984); Jalabert and Pastawski (2001); Gorin et al. (2006) and out-of-time ordered correlators (OTOCs)Sekino and Susskind (2008); Maldacena et al. (2016). Out of these many tools, Krylov Complexity — which can be defined for both states and operators — and OTOCs have emerged as the most comprehensive tests and have become the primary norm for diagnosing quantum chaos in many body systems Parker et al. (2019); Rabinovici et al. (2022); Nandy et al. (2024); Baggioli et al. (2024); Chapman et al. (2024); Jha and Roy (2024); Alishahiha et al. (2024); Swingle (2018); Lin and Motrunich (2018); García-Mata et al. (2023); Schuckert and Knap (2019).
While OTOCs have become the leading test in classifying quantum chaos and scrambling, their calculation is cumbersome and the numerical analysis of extracting Lyapunov exponents from OTOCs requires carefully fitting an exponential in a narrow parameter window. On the other hand, we have diagnostic tools that are based on Krylov subspace methods which do not require a fitting process as delicate as the OTOCs; however, due to the numerical instabilities in the Lanczos algorithm they require either a full or partial re-orthogonalization procedures limiting the system sizes that can be reached with these methods Sánchez-Garrido (2024). Here, we use both of these techniques to give a complete picture that encompasses large system sizes and a concrete diagnostic of quantum chaos and scrambling.
A summary of our results can be seen in Fig. 1, where a simple model inspired by holography — the Hyperbolic Ising model — satisfies multiple criteria of quantum chaos and fast scrambling. We evaluate the scrambling time $t_{s}$ and find a logarithmic lightcone, with $t_{s}\sim\log(N)$, where $N$ is the number of spins. The Lyapunov exponents obtained from OTOCs at finite temperatures decay as $a/\beta$, where $\beta=1/T$ is the inverse temperature and $a$ is a constant. Most importantly, the decay of these Lyapunov exponents for a certain set of parameters saturates the Maldacena-Shenker-Stanford (MSS) bound on chaos Maldacena et al. (2016). We also study the Krylov state complexity (K-complexity), which exhibits the expected ramp-peak-plateau behavior Jha and Roy (2024); Alishahiha et al. (2024), the Krylov Entropy (K-entropy) which rises exponentially at early times, and the Krylov operator complexity which shows the expected exponential-linear-saturation behavior Parker et al. (2019); Rabinovici et al. (2022); Nandy et al. (2024); all of these are indicative of fast scrambling and chaos.
Figure 1: a) Left: Tessellation of the Poincaré disk. The red line displays the line we choose to obtain the $AdS_{2}$ Ising model; the spins are located at the vertices. Right: Site-dependent nearest neighbor couplings applied to a flat spin chain to capture the curved background. Different colors on the line plot show how our site-dependent coupling changes according to different $l_{\rm max}$ values. b,c) Checks for scrambling and chaos in the hyperbolic Ising model, (b) using tensor networks for OTOCs, and (c) using Krylov subspace methods.
These tests provide strong evidence that this model exhibits quantum chaos and can be classified as a fast scrambler, giving us one of the few examples where a fast scrambling behavior has been seen outside of the SYK model and the only example with computational resources needed for its classical and quantum simulations similar to the mixed field Ising model. This discovery opens new avenues of research where we can use locally varying interactions to control different behaviors of scrambling and information propagation in quantum systems and test quantum chaos and scrambling without incurring the potentially heavy computational costs of an all-to-all connected model.
II Hyperbolic Ising Model
Our model of interest, the Hyperbolic Ising model is a mixed field Ising (MFI) model formulated on a one-dimensional hyperbolic space. In Fig. 1 we show how one can visualize this model as a line between two boundary points in $AdS_{3}$ and interpret the effects of the background curvature as site-dependent couplings.
The Hamiltonian that describes this Ising chain can be given as Asaduzzaman et al. (2024b); Ueda et al. (2011, 2010a, 2010b); Brower et al. (2022),
$\hat{H}=$ $-J\sum_{i}\left(\frac{\eta_{i}+\eta_{i+1}}{2}\right)\sigma_{i}^{z}\sigma_{i+1}^{z}$ (1)
$+h\sum_{i}\eta_{i}\sigma_{i}^{x}+m\sum_{i}\eta_{i}\sigma_{i}^{z},$
$\text{with}\hskip 5.69046pt\eta_{i}=\cosh({\rho_{i}})$
where $\sigma^{p}_{i}$ is a local Pauli operator at site $i$ with $p\in\{x,y,z\}$. The local site-dependent coupling terms for the Ising chain, $\eta_{i}=\cosh({\rho_{i}})\sim\sqrt{g}$ arise from the Euclidean metric of $\mathrm{AdS}_{2}$,
$ds^{2}$ $=\;\ell^{2}(\cosh^{2}(\rho)dt^{2}+d\rho^{2})$ (2)
To obtain the couplings in Eq. 1 we set $\ell=1.0$ and discretize radius of curvature $\rho$ as follows,
$\rho_{i}$ $=\;-l_{\rm max}+i\frac{2l_{\rm max}}{N-1}$ (3)
here $N$ is the size of the spin chain and $l_{\rm max}$ controls the curvature of the underlying curved space. One can easily recover the flat mixed field Ising model by taking the $l_{\rm max}\to 0$ limit. This model exhibits a phase transition at $J/h=1.0$ for a small value of the parameter $m$ similar to the flat Ising model Asaduzzaman et al. (2024b).
In the rest of the paper, we will center our discussion around this critical point where the mixed field Ising model is known to be chaotic Bañuls et al. (2011) and show that the inclusion of the curved background enhances this chaotic behavior and introduces fast scrambling.
Figure 2: Scrambling time $t_{s}$ as a function of site index $i$ for $N=37$, $l_{\rm max}=(1.0,...5.0)$ at infinite temperature showing the lightcone changing from linear to logarithmic as a function of $l_{\rm max}$.
Figure 3: a) $O$ for $N=13$, $J/h=1.0$, $m=0.05$, and $l_{\rm max}=4.0$ black dashed lines show the fitting range we used in extracting the Lyapunov exponents. b) The Lyapunov exponent $\lambda_{\beta}$; for $l_{\rm max}=(1.0,2.0,4.0)$. Notice that $\lambda_{\beta}$ starts saturating the MSS bound as the background curvature is increased. Error bounds are obtained from the confidence intervals of the fitting.
In an earlier work Asaduzzaman et al. (2024b), we investigated information propagation and scrambling in this model in the infinite temperature limit, and showed that for a suitable subset of the parameters the lightcone obtained from infinite temperature OTOCs is logarithmic such that the saturation time $t_{s}\sim\log(i)$ where $i$ is the site index which implies that $t_{s}\sim\log(N)$ for a system of $N$ spins. This logarithmic spreading can be seen in Fig. 2 where lines show fits to a logarithmic function, and it is one of the tell-tale signs of fast scrambling Sekino and Susskind (2008). Surprisingly, this simple model with local interactions captures this behavior quite easily due to its unique site-dependent couplings arising from the curved background contrary to the usual expectations that require all-to-all or long-range interactions to achieve such behavior.
III Observables for Chaos and Fast Scrambling
Motivated by the observations of logarithmic lightcones in this model, we will provide further evidence that this model is indeed chaotic and can be classified as a fast-scrambler by investigating finite temperature OTOCs and various complexity measures defined using Krylov subspace methods.
III.1 OTOCs at Finite Temperature
OTOCs are one of the most heavily used diagnostic tools for scrambling and chaos. They are obtained from the double commutator of operators $W(t)=e^{iHt}We^{-iHt}$ and $V$ and can be given as follows.
[…]
For a chaotic system, OTOCs are expected to decay as $F_{\beta}(t)\sim e^{-\lambda_{\beta}t}$ for times before scrambling time, where $\lambda_{\beta}$ is the Lyapunov exponent and can be thought of as the quantum counterpart of the classical Lyapunov exponent.
For systems at finite temperatures, it is common to employ a regularized definition for the OTOCs due to divergences Maldacena et al. (2016). This regularized form for the OTOCs can be defined as
[…]
where $\alpha^{4}=Z^{-1}e^{-\beta H}$.
As discussed earlier, we will choose $J/h=1.0$, and $m=0.05$ such that the model is near the chaotic region of the mixed field Ising model and choose $W_{i}$, and $V_{j}$ as $\sigma_{\frac{N+1}{2}}^{z}$ and $\sigma_{1}^{z}$ respectively and obtain the finite temperature states that go into this calculation by purification techniquesFeiguin and White (2005); Alhambra and Cirac (2021) developed for Matrix Product States (MPS) [^2] For tensor network simulations we used the ITensor Library Fishman et al. (2020); details of our simulation procedure can be found in Appendix. B. White and Feiguin (2004); White (1993); Schollwöck (2005); Vidal (2003); Verstraete et al. (2004); Vidal (2004).
After the desired thermal state is obtained it is then time evolved while measuring $O=1-\tilde{F_{ij}}(t)$ at each time-step. In Fig. 3 we show our results for the OTOCs at $l_{\rm max}=4.0$ at inverse temperatures ranging from $\beta=[0.5,...6.0]$. To extract the Lyapunov exponents we fit $\log(\frac{dO}{dt})$ to $f(x)=at+b$; this approach yields better fits with better $R^{2}$ results compared to exponential fits of the original function.
Our results presented in Fig. 3 (a) confirm the expected exponential rise of the OTOCs with the two dashed lines indicating the fitting range. This range is chosen such that for all the values of $\beta$ the linear fits result in a $R^{2}\geq 0.96$.
Fig. 3 (b) shows the decay of the $\lambda_{\beta}$ as $a/\beta$ where $a$ is some constant. Most importantly we clearly see that as the background curvature is increased the Lyapunov exponents start saturating the Maldacena-Shenker-Stanford (MSS) bound on chaos which is the most famous criterion for fast scrambling.
Figure 4: a) $C_{K}$ using states for $N=11,J/h=1.0,m=0.05$ at different $l_{\rm max}$ values showing the ramp-peak-plateau behavior. b) $S_{K}$ for the same set of parameters showing the expected exponential behavior. c) Lanczos coefficients as a function of the state index $n$ which shows the expected linear dependence on $n$.
Although our results show agreement with the MSS bound we note that this result is highly dependent on choosing the right fitting window. The dependence of $1/\beta$ is less sensitive in changes to this window but the coefficient of the fit is and thus can affect the temperature range where the MSS bound is satisfied. We also note that due to the Trotterized evolution used in both obtaining the thermal state and the time-evolved states, we are limited to moderate values of $\beta$ that can be simulated reliably due to the accumulation of Trotter errors.
Nevertheless, combining the $1/\beta$ decay of the Lyapunov exponents with the fact that we get a logarithmic lightcone for the same sets of parameters provides compelling evidence of fast scrambling.
III.2 Krylov Observables
Now we switch gears and discuss our observables obtained using Krylov subspace methods. The main advantage of these observables is that they do not require a delicate fitting scheme like OTOCs do to diagnose chaos.
The two quantities of interest are the state and operator version of the K-complexity $C_{K}$ and Krylov Entropy $S_{K}(t)$ calculated using states. These observables obtained from the Krylov basis wavefunctions exhibit different characteristics for integrable and chaotic systems allowing them to be used as a diagnostic tool for quantum chaos Parker et al. (2019); Rabinovici et al. (2022); Nandy et al. (2024); Baggioli et al. (2024); Chapman et al. (2024); Jha and Roy (2024). The Krylov complexity $C_{K}(t)$ and Krylov Entropy $S_{K}(t)$ for states can be defined as,
$C_{K}(t)$ $=\;\sum_{n=0}^{\mathbb{K}-1}n\lvert\phi_{n}(t)\rvert^{2}$ (6)
$S_{K}(t)$ $=\;-\sum_{n=0}^{\mathbb{K}-1}\lvert\phi_{n}(t)\rvert^{2}\ln{\left(\lvert\phi_{n}(t)\rvert^{2}\right)}$ ) (7)
where $\phi_{n}(t)$ are the Krylov basis wavefunctions obtained via Lanczos algorithm Lanczos (1950) with $\mathbb{K}=dim(K)$, is the dimension of the Krylov subspace $K$.
The dynamics in the Krylov subspace are fully determined by the time evolution of Krylov wavefunctions and due to the tridiagonal structure of the Lanczos coefficients time evolution in the Krylov subspace simplifies to the following recursion relation Sánchez-Garrido (2024).
$\dot{\phi}_{n}(t)=b_{n}\phi_{n-1}(t)-b_{n+1}\phi_{n+1}(t)$ (8)
Here $\{(a_{n},b_{n})\}_{n=0}^{\mathbb{K}-1}$ are the Lanczos coefficients obtained via the Lanczos algorithm Lanczos (1950) and the equations can be solved with the initial condition $\phi_{n}(t=0)=\delta_{n0}$.
Furthermore, these recursion relations map the time evolution in the full space to a single particle hopping model in the one-dimensional Krylov chain making long-time simulations tractable, especially for time scales where our tensor network methods break down. More details about the Lanczos algorithm and our implementation can be found in the Appendix. A
This brings us to our first set of observables that signals a chaotic system in the context of Krylov methods. We start the Lanczos algorithm with the state $\bigotimes_{i=1}^{N}\ket{+^{y}}$, where |$\ket{+^{y}}=\frac{1}{\sqrt{2}}\left(\ket{0}^{y}+i\ket{1}^{y}\right)$. We chose this initial seed choice because in the flat Ising model this state exhibited a clear peak in $C_{K}(t)$ in the chaotic regime of the model Alishahiha et al. (2024). [^3] The choice of the initial seed is an essential part of this algorithm and it can have an effect on the results obtained for $C_{K}(t)$ and $S_{K}(t)$. To alleviate this dependence on the initial seed new algorithms that sample a wide array of initial seeds have been proposed Craps et al. (2025). Using this initial choice we carry out the subspace generation and time evolution and investigate $C_{K}(t)$ and $S_{K}(t)$.
There are three key indicators of quantum chaos obtained using states. First, $C_{K}(t)$ is expected to exhibit ramp-peak-plateau behavior in time. Next, $S_{K}(t)$ should exhibit an exponential behavior at early timescales. Finally, the Lanczos coefficients $b_{n}$ should have a linear relationship with $n$, where $10\leq n\leq\mathbb{K}-$ 1 Parker et al. (2019); Rabinovici et al. (2022); Nandy et al. (2024); Baggioli et al. (2024); Chapman et al. (2024); Jha and Roy (2024).
Our results confirm the expectations described above and show the system becomes more chaotic as the background curvature increases. In Fig. 4 (a) we show the $C_{K}(t)$ where the ramp-peak-plateau behavior is clearly visible and the peak becomes more prominent as we increase $l_{\rm max}$ in line with the expectation that the model becomes more chaotic as $l_{\rm max}$ is increased. In Fig. 4 (b) we show $S_{K}(t)$ which shows the expected exponential rise, the lines show fits to a function of form $f(t)=a(e^{c}-d)$ and the factor $c$ increases with $l_{\rm max}$. Finally, in Fig. 4 (c), we show the Lanczos coefficient $b_{n}$ vs $n$ which shows the linear dependence on $n$, providing us with three more tests signaling chaos.
Figure 5: Operator Complexity for $N=6,J/h=1.0,m=0.05$,. Inset shows the change from exponential to linear behavior around the scrambling times.
The next set of observables are based on calculations using operators instead of states. Lanczos algorithm can be generalized to operators in straightforward way using the channel-state duality Jiang et al. (2013). In this case, the Hamiltonian becomes a Liouvillian that acts on the operators, and the usual inner product $\langle p\lvert q\rangle$ of quantum states in Lanczos Algorithm is changed to $\left(A\lvert B\right):=\frac{1}{D}Tr\left[A^{\dagger}B\right]$. We choose $\sigma^{z}_{1}$ as our initial seed operator for the Lanczos algorithm and carry out the subspace generation and time-evolution.
The $C_{K}(t)$ obtained using the operators is closely related to operator growth in the full model and hence can be used to identify scrambling behavior and chaos Nandy et al. (2024). For a fast scrambler, operator complexity is expected to grow exponentially at early timescales followed by a linear increase and finally, they saturate/oscillate about some fixed average value related to the dimension of the Krylov subspace Barbón et al. (2019).
Fig. 5 confirms this expected behavior for a chaotic system with the inset showing the early time dynamics of the complexity which is fitted to an exponential followed by a linear function. Confirming our discovery of scrambling in the Hyperbolic Ising model.
IV Conclusions and Outlook
In this paper, we investigated chaos and scrambling in the Hyperbolic Ising model which is a nearest neighbor mixed field Ising model with site-dependent couplings using tensor network and Krylov subspace methods and showed that the observables obtained from both of these independent techniques show signatures of fast scrambling and chaos. More specifically we had three main observables with such a signature.
Using tensor networks we showed that, for a suitable subset of the parameter space we observed a logarithmic lightcone with scrambling time $t_{s}\sim log(N)$. Finite temperature OTOCs showed the expected exponential rise at early times similar to the operator complexity, with Lyapunov exponents obtained from these exponentials decaying as $a/\beta$ while saturating the MSS bound for a certain set of parameters.
With the Krylov subspace methods, we showed that the observables obtained using states exhibit the following signatures of chaos. K-Complexity demonstrated the expected ramp-peak-plateau behavior with the peak scaling according to the background curvature $l_{\rm max}$. Next, K-Entropy showed the expected exponential rise at early times. Finally, we showed that Lanczos coefficients $b_{n}$s depend linearly on $n$ which is another criterion of chaos.
Operator complexity showed an exponential rise similar to the OTOCs at early times followed by a linear increase, and saturation around a maximum at late times.
Combining these results, we see robust evidence that this model can be classified as a fast scrambler. This provides us with a unique model in the class of fast scramblers where with only nearest-neighbor interactions and site-dependent couplings fast scrambling behavior can be achieved. In addition, in Fig. 2 we showed that it is possible to switch between different scrambling behaviors with ease in this model ranging from the slowest linear spreading to the fastest logarithmic spreading and anything in between by adjusting the background curvature and nearest neighbor coupling $J$.
This opens up new possibilities for studying scrambling and chaos on classical and quantum computers in detail, where we now have access to a model where by changing just two parameters we can have complete control of different kinds of scrambling behavior with modest requirements needed for its simulations compared to all-to-all connected models making it easily accessible for future studies.
In addition to the fast scrambling behavior, simulations of models that exhibit AdS/CFT correspondence are interesting in and of themselves due to the close connection between holographic models and error correcting codes Pastawski et al. (2015); Zhang et al. (2024); Chandrasekaran and Levine (2022); Wang (2024); Almheiri et al. (2015); having an efficient way of simulating these models on quantum computers could benefit the study of such codes.
Acknowledgments
We thank Raghav Govind Jha and Bharath Sambasivam for valuable discussions. AFK and GCT acknowledge financial support from the National Science Foundation under award No. 2325080: PIF: Software-Tailored Architecture for Quantum Co-Design. AS is supported by U.S. Department of Energy grant DE-SC0019139. We acknowledge the computing resources provided by North Carolina State University High Performance Computing Services Core Facility (RRID:${\rm SCR\_}022168$).
Appendix A Krylov Subspace Methods
Krylov methods are one of the widely used numerical techniques for dealing with large matrices, particularly obtaining their characteristic equations. However, for quantum mechanical systems, such methods have proven to be useful in studying time evolution of quantum states/operators without exactly diagonalizing the Hamiltonian for a system. The method is based on the Lanczos algorithm which can be summarized as follows Sánchez-Garrido (2024); Lanczos (1950),
  1. 1.
    Set |$\ket{K_{0}}=\ket{v_{0}}$, where |$\ket{v_{0}}$ is the initial state you start with.
  2. 2.
    $a_{0}=\langle K_{0}\lvert H\rvert K_{0}\rangle$
  3. 3.
    Loop over $n$, where $n\geq 1$
    1. (a)
      |$\ket{A_{n}}=H\ket{K_{n-1}}$
    2. (b)
      FRO : |$\ket{A_{n}}\to\ket{A_{n}}-\sum_{m=0}^{n-1}\langle K_{m}\lvert A_{n}\rangle\ket{K_{m}}$
    3. (c)
      $b_{n}=\sqrt{\langle A_{n}\lvert A_{n}\rangle}$. If $b_{n}=0$, exit the loop else continue.
    4. (d)
      |$\ket{K_{n}}=\frac{1}{b_{n}}\ket{A_{n}}$
    5. (e)
      $a_{n}=\langle K_{n}\lvert H\rvert K_{n}\rangle$
This algorithm has numerical instabilities due to the round-off errors and it requires either a full or a partial re-orthogonalization to be carried out to ensure the orthogonality between Krylov vectors. This procedure is basically a Gram-Schmidt process applied to Krylov vectors/operators after every iteration. Above algorithm can also be used for operators Sánchez-Garrido (2024) with redefining the inner products accordingly. Once the Lancozs coefficients are obtained using this algorithm, we can write the recursion relations and solve them.
To obtain the time evolution of these vectors, recursion relations in Eq. LABEL:eq:rec_st8 can be written in a vector-matrix linear ordinary differential equation form given in Eq. 9.
$\frac{d\vec{\phi}}{dt}=M\vec{\phi}$ (9)
where $\vec{\phi}=(\phi_{0},\phi_{1},\ldots,\phi_{\mathbb{K}-1})^{T}$ and $M$ is a tridiagonal matrix which has a different form for states $M_{s}$ and operators $M_{o}$.
These matrices are used for obtaining state and operator complexities which are given by Eq. 10 and Eq. 11 respectively.
$M_{s}=-i\begin{bmatrix}a_{0}&b_{1}&0&0&\ldots\\ b_{1}&a_{1}&b_{2}&0&\ldots\\ 0&b_{2}&a_{2}&b_{3}&\ldots\\ \vdots&\ddots&\ddots&\ddots&b_{\mathbb{K}-1}\\ 0&0&\ldots&b_{\mathbb{K}-1}&a_{\mathbb{K}-1}\par\end{bmatrix}$ (10)
$M_{o}=\begin{bmatrix}0&-b_{1}&0&0&\ldots\\ b_{1}&0&-b_{2}&0&\ldots\\ 0&b_{2}&0&-b_{3}&\ldots\\ \vdots&\ddots&\ddots&\ddots&-b_{\mathbb{K}-1}\\ 0&0&\ldots&b_{\mathbb{K}-1}&0\end{bmatrix}$ (11)
The solution to the differential equation in Eq. 9 can be given as $\vec{\phi}(t)=e^{Mt}\vec{\phi}(0)$ where $\vec{\phi}(0)=(1,0,0,\ldots,0)^{T}$ and is also expressed by the following recursion relation.
$\dot{\phi}_{n}(t)=b_{n}\phi_{n-1}(t)-b_{n+1}\phi_{n+1}(t)$ (12)
A fourth-order Runge Kutta (RK4) method is then used to solve this recursion relation numerically.
Figure 6: Lightcone at $N=13,J/h=1.0,m=0.05,\beta=0.25,l_{\rm max}=0.05$
Figure 7: Lightcone at $N=13,J/h=1.0,m=0.05,\beta=0.25,l_{\rm max}=3.0$
Appendix B Details on Thermal State Preparation
The purification procedure used to obtain the thermal states can be quickly summarized as follows.
We start with an infinite temperature Thermofield double state |$\ket{\rm{TFD}(\infty)}$ which can be initialized as
|$\ket{\rm{TFD}(\infty)}=\mathbf{1}/\sqrt{2}$ (13)
into a Matrix Product Operator(MPO) and evolved under imaginary time evolution with Time-Evolving Block Decimation(TEBD) techniques (using $e^{-\beta H}$, where $H$ is the Hyperbolic Ising Hamiltonian).
If one continues the time evolution indefinitely this algorithm eventually reaches the ground state of the model. However, stopping at a desired $\beta$ allows us to obtain the thermal states of the Hyperbolic Ising model in terms of an MPO.
With access to the thermal states of the model, we can calculate the OTOCs constructed using two local operators $W_{i}(t)$ and $V_{j}$ where $W(t)=e^{iHt}W(0)e^{-iHt}$, combined into the double commutator.
$C(t)=\langle||[W_{i}(t),V_{j}]||^{2}\rangle=2(1-\mathrm{Re}[F_{ij}(t)]),$ (14)
[…]
$F_{ij}(t)=\langle W_{i}(t)^{\dagger}V_{j}(0)^{\dagger}W_{i}(t)V_{j}(0)\rangle.$ (15)
In our calculations, we take $W(t)=\sigma^{z}(t)$, $V=\sigma^{z}$ and fix the position of the $W(t)$ operator at the center of the lattice chain which is the site $(N+1)/2$. we then place the operator $V$ at different lattice sites $i$ and measure $F_{ij}(t)$
For constructing the lightcone we need to calculate the OTOC at all sites of the spin chain which leads to the following propagation patterns in Fig. 67. As can be seen from these two plots, as the background curvature increases, the lightcone becomes warped. In Asaduzzaman et al. (2024b) we showed in detail that depending on the curvature $l_{\rm max}$ and the nearest-neighbor coupling strength $J$ it is possible to achieve different kinds of scrambling behavior ranging from linear to logarithmic using OTOCs calculated at infinite temperature.

### v2

@@
-Fast Scrambling in Hyperbolic Ising Model
+Fast Scrambling in the Hyperbolic Ising Model
 Abstract
@@
 Moreover, the modest resources needed to simulate this model make it an ideal test-bed for studying scrambling and chaos on quantum computers.
-I Introduction
-Quantum many-body chaos and scrambling have attracted a great deal of attention in recent years, from more theoretical endeavors like studying the scrambling dynamics of black holes using their dual descriptions Kobrin et al.
+Introduction— Quantum many-body chaos and scrambling have attracted a great deal of attention in recent years, from more theoretical endeavors like studying the scrambling dynamics of black holes using their dual descriptions Kobrin et al.
 (2021); Fu and Sachdev (2016); Asaduzzaman et al.
@@
 (2023); Schuster et al.
-(2022); Cotler et al.
+(2022a); Cotler et al.
 (2017); Joshi et al.
-(2022) to more practical applications like investigating scrambling in chemical reactions to quantum cryptography Chamon et al.
-(2024, 2022) and many more.
+(2022) to more practical applications like investigating scrambling in chemical reactions to quantum cryptography Zhang et al.
+(2024); Chamon et al.
+(2024) and many moreLandsman et al.
+(2019); Schuster et al.
+(2022b); Blok et al.
+(2021).
 In almost all of these studies, the Sachdev-Ye-Kitaev (SYK) model has been the center of attention as the quintessential example of a model that exhibits fast scrambling and quantum chaos Sachdev and Ye (1993); noa (a, b); Maldacena and Stanford (2016).
@@
 (2020); Yin and Lucas (2020); Bentsen et al.
-(2019).
-A wide variety of tests have been developed to diagnose and classify quantum chaos, including spectral form factors Kos et al.
+(2019); Chen and Zhou (2019).
+A wide variety of tests have been developed to diagnose and classify quantum chaos,
+including spectral form factors Kos et al.
 (2018); Liu (2018); Bertini et al.
@@
 (2006) and out-of-time ordered correlators (OTOCs)Sekino and Susskind (2008); Maldacena et al.
-(2016).
+(2016); Swingle et al.
+(2016); Xu and Swingle (2020); Swingle (2018a).
 Out of these many tools, Krylov Complexity — which can be defined for both states and operators — and OTOCs have emerged as the most comprehensive tests and have become the primary norm for diagnosing quantum chaos in many body systems Parker et al.
@@
 (2024); Jha and Roy (2024); Alishahiha et al.
-(2024); Swingle (2018); Lin and Motrunich (2018); García-Mata et al.
-(2023); Schuckert and Knap (2019).
+(2024); Swingle (2018b); Lin and Motrunich (2018); García-Mata et al.
+(2023); Schuckert and Knap (2019); Balasubramanian et al.
+(2025).
 While OTOCs have become the leading test in classifying quantum chaos and scrambling, their calculation is cumbersome and the numerical analysis of extracting Lyapunov exponents from OTOCs requires carefully fitting an exponential in a narrow parameter window.
@@
 We also study the Krylov state complexity (K-complexity), which exhibits the expected ramp-peak-plateau behavior Jha and Roy (2024); Alishahiha et al.
-(2024), the Krylov Entropy (K-entropy) which rises exponentially at early times, and the Krylov operator complexity which shows the expected exponential-linear-saturation behavior Parker et al.
+(2024), the Krylov Entropy (K-entropy), which rises exponentially at early times, and the Krylov operator complexity, which shows the expected exponential-linear-saturation behavior Parker et al.
 (2019); Rabinovici et al.
@@
 This discovery opens new avenues of research where we can use locally varying interactions to control different behaviors of scrambling and information propagation in quantum systems and test quantum chaos and scrambling without incurring the potentially heavy computational costs of an all-to-all connected model.
-II Hyperbolic Ising Model
-Our model of interest, the Hyperbolic Ising model is a mixed field Ising (MFI) model formulated on a one-dimensional hyperbolic space.
+Hyperbolic Ising Model— Our model of interest, the Hyperbolic Ising model, is a mixed field Ising (MFI) model formulated on a one-dimensional hyperbolic space.
 In Fig.
-1 we show how one can visualize this model as a line between two boundary points in $AdS_{3}$ and interpret the effects of the background curvature as site-dependent couplings.
+1, we show how one can visualize this model as a line between two boundary points in $AdS_{3}$ and interpret the effects of the background curvature as site-dependent couplings.
 The Hamiltonian that describes this Ising chain can be given as Asaduzzaman et al. (2024b); Ueda et al. (2011, 2010a, 2010b); Brower et al. (2022),
@@
 This logarithmic spreading can be seen in Fig.
-2 where lines show fits to a logarithmic function, and it is one of the tell-tale signs of fast scrambling Sekino and Susskind (2008).
-Surprisingly, this simple model with local interactions captures this behavior quite easily due to its unique site-dependent couplings arising from the curved background contrary to the usual expectations that require all-to-all or long-range interactions to achieve such behavior.
-III Observables for Chaos and Fast Scrambling
-Motivated by the observations of logarithmic lightcones in this model, we will provide further evidence that this model is indeed chaotic and can be classified as a fast-scrambler by investigating finite temperature OTOCs and various complexity measures defined using Krylov subspace methods.
-III.1 OTOCs at Finite Temperature
-OTOCs are one of the most heavily used diagnostic tools for scrambling and chaos. They are obtained from the double commutator of operators $W(t)=e^{iHt}We^{-iHt}$ and $V$ and can be given as follows.
+2, where lines show fits to a logarithmic function, and it is one of the tell-tale signs of fast scrambling Sekino and Susskind (2008).
+Surprisingly, this simple model with local interactions captures this behavior quite easily due to its unique site-dependent couplings arising from the curved background, contrary to the usual expectations that require all-to-all or long-range interactions to achieve such behavior.
+Observables for Chaos and Fast Scrambling— Motivated by the observations of logarithmic lightcones in this model, we will provide further evidence that this model is indeed chaotic and can be classified as a fast-scrambler by investigating finite temperature OTOCs and various complexity measures defined using Krylov subspace methods.
+OTOCs at Finite Temperature— OTOCs are one of the most heavily used diagnostic tools for scrambling and chaos.
+They are obtained from the double commutator of operators $W(t)=e^{iHt}We^{-iHt}$ and $V$ and can be given as follows.
 […]
@@
 where $\alpha^{4}=Z^{-1}e^{-\beta H}$.
-As discussed earlier, we will choose $J/h=1.0$, and $m=0.05$ such that the model is near the chaotic region of the mixed field Ising model and choose $W_{i}$, and $V_{j}$ as $\sigma_{\frac{N+1}{2}}^{z}$ and $\sigma_{1}^{z}$ respectively and obtain the finite temperature states that go into this calculation by purification techniquesFeiguin and White (2005); Alhambra and Cirac (2021) developed for Matrix Product States (MPS) [^2] For tensor network simulations we used the ITensor Library Fishman et al.
+As discussed earlier, we will choose $J/h=1.0$, and $m=0.05$ such that the model is near the chaotic region of the mixed field Ising model and choose $W_{i}$, and $V_{j}$ as $\sigma_{\frac{N+1}{2}}^{z}$ and $\sigma_{1}^{z}$ respectively and obtain the finite temperature states that go into this calculation by purification techniquesFeiguin and White (2005a); Alhambra and Cirac (2021); Feiguin and White (2005b) developed for Matrix Product States (MPS) [^2] For tensor network simulations we used the ITensor Library Fishman et al.
 (2020); details of our simulation procedure can be found in Appendix.
@@
 (2004); Vidal (2004).
-After the desired thermal state is obtained it is then time evolved while measuring $O=1-\tilde{F_{ij}}(t)$ at each time-step.
+After the desired thermal state is obtained, it is then time evolved while measuring $O=1-\tilde{F_{ij}}(t)$ at each time-step.
 In Fig.
 3 we show our results for the OTOCs at $l_{\rm max}=4.0$ at inverse temperatures ranging from $\beta=[0.5,...6.0]$.
-To extract the Lyapunov exponents we fit $\log(\frac{dO}{dt})$ to $f(x)=at+b$; this approach yields better fits with better $R^{2}$ results compared to exponential fits of the original function.
+To extract the Lyapunov exponent,s we fit $\log(\frac{dO}{dt})$ to $f(x)=at+b$; this approach yields better fits with better $R^{2}$ results compared to exponential fits of the original function.
 Our results presented in Fig.
 3 (a) confirm the expected exponential rise of the OTOCs with the two dashed lines indicating the fitting range.
-This range is chosen such that for all the values of $\beta$ the linear fits result in a $R^{2}\geq 0.96$.
+This range is chosen such that for all the values of $\beta$, the linear fits result in a $R^{2}\geq 0.96$.
 Fig.
 3 (b) shows the decay of the $\lambda_{\beta}$ as $a/\beta$ where $a$ is some constant.
-Most importantly we clearly see that as the background curvature is increased the Lyapunov exponents start saturating the Maldacena-Shenker-Stanford (MSS) bound on chaos which is the most famous criterion for fast scrambling.
+Most importantly, we clearly see that as the background curvature is increased, the Lyapunov exponents start saturating the Maldacena-Shenker-Stanford (MSS) bound on chaos, which is the most famous criterion for fast scrambling.
 Figure 4: a) $C_{K}$ using states for $N=11,J/h=1.0,m=0.05$ at different $l_{\rm max}$ values showing the ramp-peak-plateau behavior.
 b) $S_{K}$ for the same set of parameters showing the expected exponential behavior.
-c) Lanczos coefficients as a function of the state index $n$ which shows the expected linear dependence on $n$.
-Although our results show agreement with the MSS bound we note that this result is highly dependent on choosing the right fitting window.
-The dependence of $1/\beta$ is less sensitive in changes to this window but the coefficient of the fit is and thus can affect the temperature range where the MSS bound is satisfied.
+c) Lanczos coefficients $b_{n}$ as a function of the state index $n$, which displays the Lanczos descent.
+Although our results show agreement with the MSS bound, we note that this result is highly dependent on choosing the right fitting window.
+The dependence of $1/\beta$ is less sensitive in changes to this window,w but the coefficient of the fit is and thus can affect the temperature range where the MSS bound is satisfied.
 We also note that due to the Trotterized evolution used in both obtaining the thermal state and the time-evolved states, we are limited to moderate values of $\beta$ that can be simulated reliably due to the accumulation of Trotter errors.
 Nevertheless, combining the $1/\beta$ decay of the Lyapunov exponents with the fact that we get a logarithmic lightcone for the same sets of parameters provides compelling evidence of fast scrambling.
-III.2 Krylov Observables
-Now we switch gears and discuss our observables obtained using Krylov subspace methods.
+Krylov Observables— Now we switch gears and discuss our observables obtained using Krylov subspace methods.
 The main advantage of these observables is that they do not require a delicate fitting scheme like OTOCs do to diagnose chaos.
@@
 A
-This brings us to our first set of observables that signals a chaotic system in the context of Krylov methods.
+This brings us to our first set of observables that signal a chaotic system in the context of Krylov methods.
 We start the Lanczos algorithm with the state $\bigotimes_{i=1}^{N}\ket{+^{y}}$, where |$\ket{+^{y}}=\frac{1}{\sqrt{2}}\left(\ket{0}^{y}+i\ket{1}^{y}\right)$.
-We chose this initial seed choice because in the flat Ising model this state exhibited a clear peak in $C_{K}(t)$ in the chaotic regime of the model Alishahiha et al.
+We chose this initial seed choice because in the flat Ising model, this state exhibited a clear peak in $C_{K}(t)$ in the chaotic regime of the model Alishahiha et al.
 (2024).
-[^3] The choice of the initial seed is an essential part of this algorithm and it can have an effect on the results obtained for $C_{K}(t)$ and $S_{K}(t)$.
+[^3] The choice of the initial seed is an essential part of this algorithm, and it can have an effect on the results obtained for $C_{K}(t)$ and $S_{K}(t)$.
 To alleviate this dependence on the initial seed new algorithms that sample a wide array of initial seeds have been proposed Craps et al.
 (2025).
-Using this initial choice we carry out the subspace generation and time evolution and investigate $C_{K}(t)$ and $S_{K}(t)$.
-There are three key indicators of quantum chaos obtained using states.
+Using this initial choice, we carry out the subspace generation and time evolution and investigate $C_{K}(t)$ and $S_{K}(t)$.
+There are two key indicators of quantum chaos obtained using states.
 First, $C_{K}(t)$ is expected to exhibit ramp-peak-plateau behavior in time.
 Next, $S_{K}(t)$ should exhibit an exponential behavior at early timescales.
-Finally, the Lanczos coefficients $b_{n}$ should have a linear relationship with $n$, where $10\leq n\leq\mathbb{K}-$ 1 Parker et al.
+Parker et al.
 (2019); Rabinovici et al.
@@
 (2024); Jha and Roy (2024).
-Our results confirm the expectations described above and show the system becomes more chaotic as the background curvature increases.
+Our results confirm the expectations described above and show that the system becomes more chaotic as the background curvature increases.
 In Fig.
@@
 4 (b) we show $S_{K}(t)$ which shows the expected exponential rise, the lines show fits to a function of form $f(t)=a(e^{c}-d)$ and the factor $c$ increases with $l_{\rm max}$.
-Finally, in Fig.
-4 (c), we show the Lanczos coefficient $b_{n}$ vs $n$ which shows the linear dependence on $n$, providing us with three more tests signaling chaos.
-Figure 5: Operator Complexity for $N=6,J/h=1.0,m=0.05$,. Inset shows the change from exponential to linear behavior around the scrambling times.
+Figure 5: Operator Complexity for $N=6,J/h=1.0,m=0.05$,. The inset shows the change from exponential to linear behavior around the scrambling times.
 The next set of observables are based on calculations using operators instead of states.
@@
 (2024).
-For a fast scrambler, operator complexity is expected to grow exponentially at early timescales followed by a linear increase and finally, they saturate/oscillate about some fixed average value related to the dimension of the Krylov subspace Barbón et al.
+For a fast scrambler, operator complexity is expected to grow exponentially at early timescales followed by a linear increase and finally, they saturate/oscillate about some fixed average value related to the dimension of the Krylov subspace.
+And the Lanczos coefficients $b_{n}$ should have a linear relationship with $n$, where $10\leq n\leq\mathbb{K}-$ 1.
+This shape that the Lanczos coefficients display is generally referred to as the Lanczos descent Barbón et al.
 (2019).
-Fig.
-5 confirms this expected behavior for a chaotic system with the inset showing the early time dynamics of the complexity which is fitted to an exponential followed by a linear function.
+In Fig.
+4 (c), we show the Lanczos coefficient $b_{n}$ vs $n$, which shows the Lanczos decent behavior, and Fig.
+5 confirms this expected behavior for a chaotic system, with the inset showing the early time dynamics of the complexity, which is fitted to an exponential function followed by a linear function.
 Confirming our discovery of scrambling in the Hyperbolic Ising model.
-IV Conclusions and Outlook
-In this paper, we investigated chaos and scrambling in the Hyperbolic Ising model which is a nearest neighbor mixed field Ising model with site-dependent couplings using tensor network and Krylov subspace methods and showed that the observables obtained from both of these independent techniques show signatures of fast scrambling and chaos.
+Conclusions and Outlook— In this paper, we investigated chaos and scrambling in the Hyperbolic Ising model which is a nearest neighbor mixed field Ising model with site-dependent couplings using tensor network and Krylov subspace methods and showed that the observables obtained from both of these independent techniques show signatures of fast scrambling and chaos.
 More specifically we had three main observables with such a signature.
@@
 (2015); having an efficient way of simulating these models on quantum computers could benefit the study of such codes.
-Acknowledgments
-We thank Raghav Govind Jha and Bharath Sambasivam for valuable discussions.
+Acknowledgments— We thank Raghav Govind Jha and Bharath Sambasivam for valuable discussions.
 AFK and GCT acknowledge financial support from the National Science Foundation under award No.
@@
 Once the Lancozs coefficients are obtained using this algorithm, we can write the recursion relations and solve them.
-To obtain the time evolution of these vectors, recursion relations in Eq. LABEL:eq:rec_st8 can be written in a vector-matrix linear ordinary differential equation form given in Eq. 9.
+To obtain the time evolution of these vectors, recursion relations in Eq. 8 can be written in a vector-matrix linear ordinary differential equation form given in Eq. 9.
 $\frac{d\vec{\phi}}{dt}=M\vec{\phi}$ (9)